Phosphorus pentachloride decomposes according to the chemical equation PCl5(g)↽−−⇀PCl3(g)+Cl2(g)?c=1.80 at 250 ∘C A 0.242 mol sample of PCl5(g) is injected into an empty 3.05 L reaction vessel held at 250 ∘C. Calculate the concentrations of PCl5(g) and PCl3(g) at equilibrium.
Answer #1

Answer

[PCl5] = 0.00322M

[PCl3] = 0.07612M

Explanation

PCl5(g) <------> PCl3(g) + Cl2(g)

Kc = [PCl3][Cl2]/[PCl5] = 1.80

Initial concentration

[PCl5] = (0.242mol/3.05mol) × 1.00L = 0.07934M

[PCl3] = 0

[PCl2] = 0

change in concentration

[PCl5] = - x

[PCl3] = + x

[PCl2] = + x

Equillibrium concentration

[PCl5] = 0.07934 - x

[PCl3] = x

[Cl2] = x

so,

x2/( 0.07934 - x) = 1.80

solving for x

x = 0.07612

Therefore, at equill8brium

[PCl5] = 0.07934 - 0.07612 = 0.00322M

[PCl3] = 0.07612M

[Cl2] = 0.07612M